# When to use Graphs in RAG: A Comprehensive Analysis for Graph Retrieval-Augmented Generation

### v1

@@ 3.2 Dataset Construction
Corpus collection. Existing datasets often derive from generic sources like Wikipedia or news articles, which, while broadly accessible, lack explicit logical connections and structured domain expertise, unable to evaluate systems that require reasoning over implicit relationships or contextual hierarchies. We address these issues by (i) integrating tightly structured domain data from NCCN medical guidelines to embed explicit hierarchies and standardized protocols, which provide dense conceptual relationships (e.g., treatment protocols linking symptoms, drugs, and outcomes) at scales exceeding typical domain corpora, and (ii) collecting loosely organized texts (pre-20th-century novels) from Gutenberg library to simulate real-world documents with implicit, non-linear narratives, ensuring the corpus reflects the complexity of unstructured knowledge while minimizing pretraining contamination. This combination ensures the corpus balances unstructured, real-world ambiguity with domain-specific hierarchies, enabling rigorous evaluation of both retrieval robustness and reasoning depth. We include more details about these datasets (Novel and Medical Datasets) in Appendix B.
@@
Logic and evidence extraction To overcome the superficial treatment of reasoning in existing benchmarks, where multi-hop queries often reduce to linear fact retrieval, our framework systematically transforms raw text into structured domain ontologies. These ontologies preserve not only entities but also their contextual relationships and hierarchical dependencies, enabling the extraction of fine-grained evidence that reflects both localized factual clusters and interconnected reasoning chains. Where prior work struggles to represent latent logical synthesis (e.g., inferring causal pathways from dispersed market factors), our evidence extraction process isolates self-contained subgraphs for basic retrieval while reconstructing multi-hop relational sequences that expose deeper inferential patterns.
@@
Question generation we generate the questions according to the complexity of the underlying evidence. Rather than treating difficulty as a function of factual scarcity or hop count, we calibrate questions by progressively integrating evidence types, from isolated subgraph fragments for retrieval tasks to global topology-aware reasoning for synthetic reasoning. This ensures that complex questions necessitate not merely aggregating discrete facts but synthesizing contextual hierarchies and domain-specific ontologies. By anchoring questions in structured evidence packages that mirror real-world knowledge interdependencies, our benchmark evaluates how models derive insights from both explicit logical frameworks and unstructured contextual ambiguity, thereby addressing the critical gap in assessing reasoning depth beyond simple retrieval.
@@
Relevance check and refinement To ensure the accuracy and practical relevance of the dataset, we implemented rigorous validation and refinement processes after initial construction. Full methodological details are provided in B, while the visualization of our datasets is in Appendix C.
@@ 3.3 Evaluation Metrics
Average $\textsc{Average Degree}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\deg(v),$ (1)
@@
To evaluate the retrieval performance of GraphRAG, we argue that an effective system should not only ensure the completeness of retrieved information (i.e., high recall) but also reduce irrelevant content (i.e., high relevance). We introduce two corresponding retrieval-quality-based metrics as detailed below.
$\textsc{Context Relevance}=\frac{r_{1}+r_{2}}{2},\quad r_{1},r_{2}\in\{0,0.5,1\},$ (3)
where $r_{1}$ and $r_{2}$ are normalized relevance ratings. Details are provided in Appendix D.
Evidence $\textsc{Evidence Recall}=\frac{1}{|\mathcal{R}|}\sum_{c\in\mathcal{R}}\mathds{1}\left(\operatorname{S}\left(c,\mathcal{C}\right)\right),$ (4)
@@
After retrieval, a GraphRAG system is expected to generate accurate answers based on the retrieved contexts. To evaluate the quality of the generation, we introduce four key metrics: 1) Lexical Overlap: Measures word-level similarity between the generated and reference answers using longest common subsequence matching. 2) Answer Accuracy: Assesses both semantic similarity and factual consistency with the reference answer. 3) Faithfulness: Evaluates whether the relevant knowledge points in a long-form answer are faithful to the given context. 4) Evidence Coverage: Measures whether the answer adequately covers all knowledge relevant to the question. We provide details of these widely used metrics in Appendix D.
@@ 4.1 Generation Accuracy (Q1)
Figure 5: Take-away results.
To address Q1, we evaluate seven representative GraphRAG frameworks on our benchmark, using tailored metrics for different question types. For Type 1 (retrieval) and Type 2 (reasoning) questions, we assess answer quality with ROUGE scores and accuracy. For Type 3 (summarization) questions, we introduce evidence coverage to measure the comprehensiveness of the generated answers. For Type 4 (creative generation) questions, we use faithfulness to assess factual consistency. Main results in Table 2 and Appendix E lead to following observations:
Table 2: Result of Generate Evaluation,include tasks with varying comlexity from different datasets.
@@
GraphRAG (local) edge2024local 49.29 26.11 50.93 24.09 64.40 75.58 39.10 55.44 35.65
GraphRAG (global) edge2024local 36.92 17.32 43.17 15.12 56.87 80.55 41.11 75.15 30.34
HippoRAG HippoRAG 52.93 26.65 38.52 11.16 48.70 85.55 38.85 71.53 38.97
HippoRAG2 gutiérrez2025hipporag2 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
LightRAG guo2024lightrag 58.62 35.72 49.07 24.16 48.85 63.05 23.80 57.28 25.01
Fast-GraphRAG fastgraphrag2024 56.95 35.90 48.55 21.12 56.41 80.82 46.18 57.19 36.99
Graph RAG RAPTOR sarthi2024raptor 49.25 23.74 38.59 11.66 47.10 82.33 38.01 70.85 35.88
@@
HippoRAG HippoRAG 56.14 20.95 55.87 13.57 59.86 62.73 64.43 69.21 65.56
HippoRAG2 gutiérrez2025hipporag2 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
LightRAG guo2024lightrag 63.32 37.19 61.32 24.98 63.14 51.16 - - -
Fast-GraphRAG fastgraphrag2024 60.93 31.04 61.73 21.37 67.88 52.07 65.93 56.07 44.73
GraphRAG RAPTOR sarthi2024raptor 54.07 17.93 53.20 11.73 58.73 78.28 - - -
@@
Obs.2 . GraphRAG excels in complex tasks: GraphRAG models show a clear advantage in complex reasoning, contextual summarization, and creative generation. This is intuitive, as these tasks require bridging the complex relations among multiple concepts, which is naturally a graph structure.
Obs.3 . GraphRAG ensures greater factual reliability in creative tasks: RAPTOR scores highest in faithfulness (70.9%) on the novel dataset, though RAG covers more evidence (40.0%)—likely because GraphRAG’s fragmented knowledge retrieval complicates broad-scope generation. This trade-off highlights GraphRAG’s strength in precision but limitations in wide-ranging synthesis.
@@ 4.2 Retrieval Performance (Q2)
To quantitatively compare the retrieval effectiveness of the two paradigms, we adopt two complementary metrics: Context Recall, which measures how completely the retrieved context covers the gold evidence, and Context Relevance, which measures the semantic alignment between the retrieved content and the input query. As shown in Table 3 and Appendix E, we have the following observtions:
Obs.4 . RAG excels at retrieving discrete facts for simple questions that do not require complex logics, achieving 83.2% Context Recall on the novel dataset (vs. HippoRAG2’s best Context Relevance). Medical dataset results confirm this pattern, suggesting relevant evidence for Level 1 questions typically resides in single passages. It is because the graph used in GraphRAG introduces several logically-relevant but redundant information in these scenarios.
@@
Table 3: Retrieval performance across tasks with varying complexity from different datasets.
@@
GraphRAG (local)edge2024local 61.04 27.30 73.03 39.09 82.02 43.13 53.55 35.07
GraphRAG (global)edge2024local 42.27 9.37 86.68 14.36 89.69 15.35 83.14 19.40
HippoRAGHippoRAG 80.44 56.34 87.91 58.75 90.95 59.46 65.51 46.64
HippoRAG2gutiérrez2025hipporag2 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
LightRAGguo2024lightrag 73.69 33.08 85.52 37.46 87.59 38.02 71.72 38.06
Fast-GraphRAGfastgraphrag2024 64.48 47.86 73.51 55.21 78.58 49.74 56.31 46.27
Graph RAG RAPTORsarthi2024raptor 62.14 54.08 67.80 61.26 75.79 63.00 58.66 58.46
@@
HippoRAG HippoRAG 87.25 52.44 83.80 42.19 83.46 49.13 81.66 45.03
HippoRAG2 gutiérrez2025hipporag2 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
LightRAG guo2024lightrag 80.32 41.27 82.91 42.79 85.71 43.11 81.34 45.17
Fast-GraphRAG fastgraphrag2024 66.82 45.86 74.93 38.80 77.27 47.58 62.99 25.15
Graph RAG RAPTOR sarthi2024raptor 85.40 69.38 89.70 53.20 88.86 58.73 72.70 52.71
@@ 4.3 Graph Complexity (Q3)
Figure 6: The illustration of graph construction.
@@
Metric GraphRAG HippoRAG2 LightRAG Fast-GraphRAG HippoRAG
@@
Obs.7 . The index graphs generated by different GraphRAG implementations demonstrate substantial structural variation. As illustrated in Figure 7, HippoRAG2 produces significantly denser graphs, with both node and edge counts exceeding those of other frameworks by considerable margins. Specifically, on the novel dataset, HippoRAG2 has an average of 2,310 edges and 523 nodes, while on the medical dataset, it has an average of 3,979 edges and 598 nodes. This enhanced graph density improves both information connectivity and coverage, ultimately contributing to superior retrieval and generation capabilities. Please refer more analysis on graph quality in Appendix E.
@@ 4.4 Efficiency (Q4)
Figure 8: Avg tokens of different GraphRAG (part 1).
Avg Tokens V-RAG GraphRAG (local) GraphRAG (global) HippoRAG2
@@
Figure 9: Avg tokens of different GraphRAG (part 2).
@@
Obs.8 . Compared to vanilla RAG, GraphRAG significantly increases prompt length due to the additional steps involved in knowledge retrieval and graph-based aggregation. Specifically, Global-GraphRAG, which incorporates a community-summarization mechanism, reaches a prompt size of up to $4\times 10^{4}$ tokens. LightRAG also produces lengthy prompts ($\approx 10^{4}$ tokens). In contrast, HippoRAG2 maintains a more compact prompt size ($\approx 10^{3}$ tokens), showing better efficiency. These results highlight that GraphRAG’s structured pipeline incurs non-trivial token overhead.
Obs.9 . As task complexity and the number of required knowledge points increase, GraphRAG’s prompt length exhibits a clear upward trend. Notably, Global-GraphRAG’s prompt size expands from 7,800 to 40,000 tokens across tasks of increasing difficulty. This excessive token accumulation often introduces redundant information, which in turn degrades context relevance during retrieval. These findings underscore a critical trade-off in GraphRAG: while it improves retrieval breadth and organization, it may also lead to inefficiencies due to prompt inflation, especially in complex tasks.
@@ A.1 Code and Leaderboard
To promote transparency and reproducibility, we have uploaded our code to GitHub at https://github.com/GraphRAG-Bench/GraphRAG-Benchmark. This repository includes the source code and scripts for data preprocessing, benchmark construction, model training, and evaluation, ensuring that researchers have full access to the resources required to reproduce and extend our work. Besides that, the leaderboard and all other related resources and analyses are collected for the community at https://graphrag-bench.github.io/.

### v2

@@ 3.2 Dataset Construction
Corpus collection. Existing datasets often derive from generic sources like Wikipedia or news articles, which, while broadly accessible, lack explicit logical connections and structured domain expertise, unable to evaluate systems that require reasoning over implicit relationships or contextual hierarchies. We address these issues by (i) integrating tightly structured domain data from NCCN medical guidelines to embed explicit hierarchies and standardized protocols, which provide dense conceptual relationships (e.g., treatment protocols linking symptoms, drugs, and outcomes) at scales exceeding typical domain corpora, and (ii) collecting loosely organized texts (pre-20th-century novels) from Gutenberg library to simulate real-world documents with implicit, non-linear narratives, ensuring the corpus reflects the complexity of unstructured knowledge while minimizing pretraining contamination. This combination ensures the corpus balances unstructured, real-world ambiguity with domain-specific hierarchies, enabling rigorous evaluation of both retrieval robustness and reasoning depth. We include more details about these datasets (Novel and Medical Datasets) in Appendix C.
@@
Logic and evidence extraction. To overcome the superficial treatment of reasoning in existing benchmarks, where multi-hop queries often reduce to linear fact retrieval, our framework systematically transforms raw text into structured domain ontologies. These ontologies preserve not only entities but also their contextual relationships and hierarchical dependencies, enabling the extraction of fine-grained evidence that reflects both localized factual clusters and interconnected reasoning chains. Where prior work struggles to represent latent logical synthesis (e.g., inferring causal pathways from dispersed market factors), our evidence extraction process isolates self-contained subgraphs for basic retrieval while reconstructing multi-hop relational sequences that expose deeper inferential patterns.
@@
Question generation. we generate the questions according to the complexity of the underlying evidence. Rather than treating difficulty as a function of factual scarcity or hop count, we calibrate questions by progressively integrating evidence types, from isolated subgraph fragments for retrieval tasks to global topology-aware reasoning for synthetic reasoning. This ensures that complex questions necessitate not merely aggregating discrete facts but synthesizing contextual hierarchies and domain-specific ontologies. By anchoring questions in structured evidence packages that mirror real-world knowledge interdependencies, our benchmark evaluates how models derive insights from both explicit logical frameworks and unstructured contextual ambiguity, thereby addressing the critical gap in assessing reasoning depth beyond simple retrieval.
@@
Relevance check and refinement. To ensure the accuracy and practical relevance of the dataset, we implemented rigorous validation and refinement processes after initial construction. Full methodological details are provided in Appendix C, while the visualization of our datasets is in Appendix E.
@@ 3.3 Evaluation Metrics
$\textsc{Average Degree}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\deg(v),$ (1)
@@
To evaluate the retrieval performance of GraphRAG, we argue that an effective system should not only ensure the completeness of retrieved information (i.e., high recall) but also reduce irrelevant content (i.e., high relevance). We introduce two corresponding retrieval-quality-based metrics: 1) Context Relevance measures how well the retrieved content aligns with the question’s intent by calculating the semantic similarity between the question and the retrieved context. 2) Evidence Recall measures retrieval completeness by assessing whether all critical components required to correctly answer the question are captured. Details are provided in Appendix F.
@@
After retrieval, a GraphRAG system is expected to generate accurate answers based on the retrieved contexts. To evaluate the quality of the generation, we introduce four key metrics: 1) Lexical Overlap: Measures word-level similarity between the generated and reference answers using longest common subsequence matching. 2) Answer Accuracy: Assesses both semantic similarity and factual consistency with the reference answer. 3) Faithfulness: Evaluates whether the relevant knowledge points in a long-form answer are faithful to the given context. 4) Evidence Coverage: Measures whether the answer adequately covers all knowledge relevant to the question. We provide details of these widely used metrics in Appendix F.
@@ 4.1 Generation Accuracy (Q1)
Figure 5: Retrieval and generation performance of RAG and GraphRAG across four different tasks.
To address Q1, we evaluate seven representative GraphRAG frameworks on our benchmark, using tailored metrics for different question types. For Type 1 (retrieval) and Type 2 (reasoning) questions, we assess answer quality with ROUGE scores and accuracy. For Type 3 (summarization) questions, we introduce evidence coverage to measure the comprehensiveness of the generated answers. For Type 4 (creative generation) questions, we use faithfulness to assess factual consistency. Main results in Table˜2 and Appendix G lead to following observations:
Table 2: Results of Generate Evaluation using GPT-4o-mini, covering tasks of varying complexity.
@@
MS-GraphRAG (Edge et al., 2024) 49.29 26.11 50.93 24.09 64.40 75.58 39.10 55.44 35.65
HippoRAG (Gutiérrez et al., 2024) 52.93 26.65 38.52 11.16 48.70 85.55 38.85 71.53 38.97
HippoRAG2 (Gutiérrez et al., 2025) 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
LightRAG (Guo et al., 2024) 58.62 35.72 49.07 24.16 48.85 63.05 23.80 57.28 25.01
Fast-GraphRAG (CircleMind-AI, 2024) 56.95 35.90 48.55 21.12 56.41 80.82 46.18 57.19 36.99
Graph RAG RAPTOR (Sarthi et al., 2024) 49.25 23.74 38.59 11.66 47.10 82.33 38.01 70.85 35.88
Lazy-GraphRAG (Darren Edge, 2024) 51.65 36.97 49.22 23.48 58.29 76.94 43.23 50.69 39.74
@@
MS-GraphRAG (Edge et al., 2024) 38.63 26.80 47.04 21.99 41.87 22.98 53.11 32.65 39.42
HippoRAG (Gutiérrez et al., 2024) 56.14 20.95 55.87 13.57 59.86 62.73 64.43 69.21 65.56
HippoRAG2 (Gutiérrez et al., 2025) 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
LightRAG (Guo et al., 2024) 63.32 37.19 61.32 24.98 63.14 51.16 67.91 78.76 51.58
Fast-GraphRAG (CircleMind-AI, 2024) 60.93 31.04 61.73 21.37 67.88 52.07 65.93 56.07 44.73
GraphRAG RAPTOR (Sarthi et al., 2024) 54.07 17.93 53.20 11.73 58.73 78.28 62.38 59.98 63.63
Lazy-GraphRAG (Darren Edge, 2024) 60.25 31.66 47.82 22.68 57.28 55.92 62.22 30.95 43.79
@@
Obs.2 . GraphRAG excels in complex tasks: GraphRAG models show a clear advantage in complex reasoning, Contextual Summarize, and creative generation. This is intuitive, as these tasks require bridging the complex relations among multiple concepts, which is naturally a graph structure.
Obs.3 . GraphRAG ensures greater factual reliability in creative tasks: RAPTOR scores highest in faithfulness (70.9%) on the novel dataset, though RAG covers more evidence (40.0%), likely because GraphRAG’s fragmented knowledge retrieval and complicates broad scope generation. This trade-off highlights GraphRAG’s strength in precision but limitations in wide-ranging synthesis.
@@ 4.2 Retrieval Performance (Q2)
To quantitatively compare the retrieval effectiveness of the two paradigms, we adopt two complementary metrics: Evidence Recall, which measures how completely the retrieved context covers the gold evidence, and Context Relevance, which measures the semantic alignment between the retrieved content and the input query. As shown in Table 3 and Appendix G, we have the following observtions:
Obs.4 . RAG excels at retrieving discrete facts for simple questions that do not require complex logics, achieving 83.2% Evidence Recall on the novel dataset (vs. HippoRAG2’s best Context Relevance). Medical dataset results confirm this pattern, suggesting relevant evidence for Level 1 questions typically resides in single passages. It is because the graph used in GraphRAG introduces several logically relevant but redundant information in these scenarios.
@@
Table 3: Results of Retrieval performance using GPT-4o-mini, covering tasks of varying complexity.
@@
MS-GraphRAG(Edge et al., 2024) 61.04 27.30 73.03 39.09 82.02 43.13 53.55 35.07
HippoRAG(Gutiérrez et al., 2024) 80.44 56.34 87.91 58.75 90.95 59.46 65.51 46.64
HippoRAG2(Gutiérrez et al., 2025) 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
LightRAG(Guo et al., 2024) 73.69 33.08 85.52 37.46 87.59 38.02 71.72 38.06
Fast-GraphRAG(CircleMind-AI, 2024) 64.48 47.86 73.51 55.21 78.58 49.74 56.31 46.27
Graph RAG RAPTOR(Sarthi et al., 2024) 62.14 54.08 67.80 61.26 75.79 63.00 58.66 58.46
Lazy-GraphRAG (Darren Edge, 2024) 59.25 30.76 57.73 42.98 77.38 43.62 55.24 31.94
@@
MS-GraphRAG (Edge et al., 2024) 38.06 05.67 61.32 04.25 59.66 05.24 66.59 02.76
HippoRAG (Gutiérrez et al., 2024) 87.25 52.44 83.80 42.19 83.46 49.13 81.66 45.03
HippoRAG2 (Gutiérrez et al., 2025) 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
LightRAG (Guo et al., 2024) 80.32 41.27 82.91 42.79 85.71 43.11 81.34 45.17
Fast-GraphRAG (CircleMind-AI, 2024) 66.82 45.86 74.93 38.80 77.27 47.58 62.99 25.15
Graph RAG RAPTOR (Sarthi et al., 2024) 85.40 69.38 89.70 53.20 88.86 58.73 72.70 52.71
Lazy-GraphRAG (Darren Edge, 2024) 74.29 19.90 78.65 17.50 78.72 21.35 83.41 15.09
@@ 4.3 Graph Complexity (Q3)
Obs.7 . The index graphs generated by different GraphRAG implementations demonstrate substantial structural variation.
Figure 6: The relational structure of different methods.
@@
Metric MS-GraphRAG HippoRAG2 LightRAG Fast-GraphRAG HippoRAG
@@
As illustrated in Figure 7, HippoRAG2 produces significantly denser graphs, with node and edge counts that far surpass other frameworks. Specifically, on the novel dataset, HippoRAG2 has an average of 2,310 edges and 523 nodes, while on the medical dataset, it has an average of 3,979 edges and 598 nodes. This enhanced graph density improves both information connectivity and coverage, ultimately contributing to superior retrieval and generation capabilities. This observation is consistent with the retrieval performance, which shows that HippoRAG2 achieves higher recall than other baseline models.
@@ 4.4 Efficiency (Q4)
Figure 8: Ave token cost of different GraphRAG (Part 1).
Avg Tokens V-RAG MS-GraphRAG(local) MS-GraphRAG(global) HippoRAG2
@@
Figure 9: Ave token cost of different GraphRAG (Part 2).
@@
Obs.8 . Compared to vanilla RAG, GraphRAG significantly increases prompt length due to the additional steps involved in knowledge retrieval and graph-based aggregation. Specifically, MS-GraphRAG(global), which incorporates a community-summarization mechanism, reaches a prompt size of up to $4\times 10^{4}$ tokens. LightRAG also produces lengthy prompts ($\approx 10^{4}$ tokens). In contrast, HippoRAG2 maintains a more compact prompt size ($\approx 10^{3}$ tokens), showing better efficiency. These results highlight that GraphRAG’s structured pipeline incurs non-trivial token overhead.
Obs.9 . As task complexity and the number of required knowledge points increase, GraphRAG’s prompt length exhibits a clear upward trend. Notably, MS-GraphRAG(global)’s prompt size expands from 7,800 to 40,000 tokens across tasks of increasing difficulty. This excessive token accumulation often introduces redundant information, which in turn degrades context relevance during retrieval. These findings underscore a critical trade-off: while GraphRAG improves retrieval breadth, it may also introduce noisy context due to prompt inflation, especially in complex tasks.
@@ 5 Conclusion
Ethics statement
The benchmark introduced in this paper is constructed from publicly available internet data. We have taken care to ensure our data collection process respects user privacy by filtering for personally identifiable information and complies with the terms of the source platforms. All models used in our evaluation are open-source, promoting transparency and reproducibility. While we have focused on the technical aspects of reasoning, we advise users to be mindful of potential societal biases that may exist in the source data and to consider the broader context of their applications.
Reproducibility statement
To ensure the reproducibility of our research, we have made our dataset, evaluation code, and detailed experimental settings publicly available. We have open-sourced the original data for our benchmark and all the code required to replicate the results presented in this paper. These resources are available in an anonymous repository at https://anonymous.4open.science/r/GraphRAG-Benchmark-CE8D/. Furthermore, we provide a detailed description of our dataset construction process in Appendix C. The specific hyperparameter settings used for all baseline models evaluated in our experiments are also fully documented in Appendix H.2.
@@ A.1 Code and Leaderboard
To promote transparency and reproducibility, we have uploaded our code to GitHub at https://github.com/GraphRAG-Bench/GraphRAG-Benchmark. This repository includes the source code and scripts for evaluation, ensuring that researchers have full access to the resources required to reproduce and extend our work. Besides that, the leaderboard and all other related resources and analyses are collected for the community at https://graphrag-bench.github.io/
Figure 10: Overview of the leaderboard, ranked by average generation performance (ACC).